OpenStudy (anonymous):

A rock is tossed into a pond creating an expanding circular ripple. If the radius of the ripple is increasing at a rate of 1.5 ft/sec, find out how fast the area is increasing when the radius is 2 feet.

OpenStudy (anonymous):

help

OpenStudy (anonymous):

dr/dt=1.5 Since A=pi*r^2/4 differentiate both sides to give rate of area increase. dA/dt=pi/4*2r*(dr/dt) plug in values for dr/dt and r

OpenStudy (anonymous):

Is it not the case that A=Pi * r^2 ?